# Shot-to-shot acquisition ultrafast electron diffraction

### v1

@@ Abstract
We demonstrate a novel shot-to-shot acquisition method for optical pump - keV electron energy probe in ultrafast scattering experiments. We integrate a phase-locked acquisition scheme at a repetition rate of 20kHz in a conventional ultrafast electron diffraction (UED) setup. We proceed to a full characterization of the noise level in different configurations and of the temporal resolution as a function of electron flux. The shot-to-shot acquisition improves the signal-to-noise ratio (SNR) by one order of magnitude in a low perturbation regime.
@@ I Introduction
Ultrafast electron diffraction, microscopy, and energy loss spectroscopy are widely used to investigate the properties of materials out-of-equilibrium using pump-probe methodsZewail (2010). Transient signals recorded by these techniques correspond to a variation in the number of scattered electrons at specific positions in the momentum, real, or energy space, which enables studies in the temporal domain of lattice dynamics Carbone et al. (2008); Chatelain et al. (2014), magnetic ordering Rubiano da Silva et al. (2018a); Berruto et al. (2018); Rubiano da Silva et al. (2018b), plasmon dynamics Piazza et al. (2015), charge density waves coherent response and melting Cheng et al. (2024), and to resolve changes in the phonon population with momentum resolutionRené De Cotret et al. (2019); Stern et al. (2018); Waldecker et al. (2017). The amplitude of transient signals increases with the photoexcitation energy density and is typically in the order of percent for light excitation in the range of mJ/cm2. Advanced material science rely on broken-symmetry ground states where a delicate microscopic balance of electron, spin, lattice and orbital degree of freedom is attained. While elastically scattered electrons provide measurements of lattice order, inelastically scattered electrons contain crucial information on the momentum-dependent interaction between collective modes van der Veen, Penfold, and Zewail . For instance, phonon populations alter thermal diffuse scattering, which has a typically total cross-section of $10^{-4}$. Other quasiparticles such as magnons are also present in the diffuse inelastic scattering but represent only $10^{-6}$ of the total electron flux Lyon et al. (2021). Time-resolved measurements of such inelastic signals present a significant challenge for current experimental setups, which must improve detection sensitivity to capture small photo-induced variations, particularly in weak perturbation regimes that maintain the integrity of the ground state.
@@
To improve the signal-to-noise ratio (SNR), a first possibility is to develop a brighter electron source ($>10^{5}$ electrons/pulse). However, this approach introduces space-charge-induced temporal broadening of the electron pulse, which significantly degrades the time resolution Siwick et al. (2002); Wang et al. (2009). Temporal compression of the electron bunch can be achieved using a synchronized radio frequency cavity Mancini et al. (2012); Chatelain et al. (2012) or a terahertz (THz) optical pulse Ehberger et al. (2019). Nevertheless, special care must be taken to avoid jitter and long-term drifts Otto et al. (2017). A second approach is to minimize the noise of the electron source, while working at low electron flux ($<10^{4}$ electron/pulse). To mitigate space charge broadening, the electron source is placed as close as possible to the sample Waldecker, Bertoni, and Ernstorfer (2015). Combined with a high repetition rate laser source and fast acquisition systemFreelon et al. (2023); Aidelsburger et al. (2010), the SNR is increased while maintaining good beam coherence. A possible limitation of this method is the sample re-equilibration time, which can become greater than the inter-pulse separation ($<1\mu$s). In this work, we describe an ultrafast electron scattering apparatus based on a source with a pulse repetition rate of 20 kHz operating at a low flux of a thousand electrons per pulse. In this limit, the quality of the acquisition system is essential to improve the SNR. We employ a direct electron detector [^1] Quadro, Dectris, which records high SNR electron diffraction patterns McMullan, Faruqi, and Henderson (2016) thanks to a combination of high detective quantum efficiency (DQE) (0.9 for direct detection, as compared to 0.2 for a typical CCD) Levin (2021) and low acquisition noise. Another advantage of such a direct detection system is its capability to acquire data at a high rate, synchronized with the laser repetition rate Duncan et al. (2023). This enables the collection of pump-probe data by chopping the pump pulse train at 10 kHz and normalizing the data using unpumped acquisition frames. This method, well-established in ultrafast optical spectroscopy due to the availability of fast light detectors, is widely used to achieve shot-noise-limited data Lang (2018). For electron scattering experiments, where the number of events per pixel per laser pulse is relatively low, this approach is feasible due to the detector’s negligibly low dark count rate. We demonstrate pump-probe data with a noise level reaching $10^{-4}$ in a few hours of acquisition time, with a time resolution in the sub-picosecond range.
@@
The paper is organized as follows. In section II, we describe the setup, the acquisition system, and the working principles of the synchronization to achieve the shot-to-shot acquisition. Section III presents the characterization of the noise level and the time resolution. In section IV, we will present UED data acquired in a realistic scenario by photo-exciting graphite samples with an 800 nm pump.
@@ II Design of the shot-to-shot UED
Figure 1: Scheme of the UED setup in EPFL composed of the optical table, the rack, and the computer. The optical table supports the pulsed laser, the optical beam path divided into pump and probe arms, the vacuum sample chamber, the electron and photodiode detector. The rack accommodates the synchronization system which is phase-locked with the laser source at 20 kHz. The data gathered by the electron and photodiode detector are sent the computer to be sorted between pumped and unpumped diffraction patterns.
@@ II.1 Optical setup and electron diffraction beamline
The ultrafast electron diffraction (UED) setup and the data acquisition scheme are schematized in Fig. 1. The current setup is a modification of the setup detailed in Ref. Mancini et al. (2012). A cryogenically cooled Ti:Sapphire laser [^2] Wyvern, KM Labs provides 50 fs pulses at a central wavelength of 780 nm at a repetition rate of 20 kHz with an output power higher than 10 W (pulse energy > 0.5 mJ). To avoid long-term drifts, the laser beam pointing is stabilized by a home-built stabilization system based on a pair of piezoelectric-actuated mirrors and CMOS cameras. The beam is divided into two arms with a beam splitter (BS1).
The pump beam goes through a mechanical delay line, to vary the delay between the probe pulse and the pump pulse. Then, is passed through a mechanical chopper, synchronized with the laser source to halve the pulse repetition rate to 10 kHz. The last mirror (BS2) sends the pump into the chamber and is used to find spatial overlap between the pump beam and the electron probe. BS2 also acts as a beam splitter, transmitting a small fraction of the pump towards a photodiode.
@@
In the probe arm, third harmonic generation (THG) is used to generate a UV light beam (260 nm) which back-illuminates a silver-coated sapphire photo-cathode. The UV probe pulse energy exceeds the work function of the silver ($P_{\text{UV}}=4.76\text{ eV}>W_{\text{Ag}}=4.26$ eV), thus generating a pulse of electrons by single-photon photoemission, with a measured quantum efficiency of $10^{-8}$ electron/photon (see Appendix A). These pulses are then accelerated up to 40 keV over a distance of 3 mm corresponding to an accelerating electric field of 13.3 MV/m. The electron beam is focused by a homemade magnetic lens consisting of around 3000 windings, each of them carrying a current of 0.62 A.
@@
After 150 mm of propagation length, the electron pulses reach the sample holder position with sub-ps pulse duration, depending on the electron flux, as described in Sec III.3. In transmission geometry, a copper sample holder hosts five standard 3 mm TEM grids. The holder is mounted on a 4-axis open-cycle cryo-manipulator that can achieve a temperature of 4.2K with liquid helium cooling. The UED setup can also work in reflection geometry to probe the surface lattice dynamics with five axes of freedom using another sample holder Pennacchio et al. (2017).
@@
The vacuum chamber has a base pressure of $10^{-8}$ mbar and accommodates a mirror that reflects the pump beams onto the sample with an angle of a few degrees off from the electron beam.
@@
The electron beam travels a distance of 491 mm after the sample position. At this distance and with a chip detection size of $38.4\times 38.4$ mm2, a reciprocal space area of $16\times 16$-2 is covered at 40 keV.
@@ II.2 High speed lock-in direct electron detection
We use this detector feature to acquire ultrafast electron diffraction by chopping the pump pulses at half the laser repetition rate (10 kHz). The two output images correspond to the excited, $I_{\text{on}}$, and nonexcited, $I_{\text{off}}$, diffraction patterns of the studied sample. Due to the high shot-to-shot correlation of the laser source, this reduces considerably the noise in the pump-probe signal, as will be discussed in Sec. III. We now describe the synchronization system in greater detail.
@@ II.3 Synchronization system
As schematized in the left panel of Fig. 1, a homemade delay generator constantly receives the 20 kHz laser signal from the laser source. This trigger is relayed by a homemade delay generator into four independently-delayed channels (A-D). Channel A is frequency-divided to 10 kHz and transmitted to the chopper controller. Channel B triggers a fast photodiode digitizer  [^3] GlazPD, Synertronic design. This signal on this channel is gated by the main detector (EnOUT) which is a replica of the one received (EnIN). This way, the Quadro detector operates as master and the photo-diode as slave, ensuring that the same pulses originating from the laser are detected. The electron detector acquisition window is controlled by channel D, with a 20 kHz repetition rate with a window of 25 $\mu$s. The acquisition window from channel D is inhibited by channel C, acting like a switch controlled by the main computer to start the acquisition without cutting off pulses. This synchronization system allows a robust acquisition of every single electron bunch at 20 kHz, but can easily be adjusted to other repetition rates although the chopping mechanism would be different for higher frequencies. From the main computer, the number of exposition windows, $n$, as well as other settings of the detector can be monitored.
The main computer accesses the two 2D datasets acquired by the electron detector, $I_{\text{A}}$ and $I_{\text{B}}$ that can be sorted into $I_{\text{on}}$ and $I_{\text{off}}$, thanks to the photodiode. The GUI can live display these data and the difference between $I_{\text{on}}-I_{\text{off}}$ allowing for an accurate time and spatial overlap.
@@ III Characterization
Given the low shot-to-shot fluctuation of the laser source, this acquisition method can greatly improve the SNR of ultrafast experiments. We characterize this enhancement by disentangling the noise sources for conventional and shot-to-shot acquisition techniques with and without the photoexcitation pulse. We also investigate the electron pulse duration of our system as a function of the flux to find the optimal compromise between electron flux and time resolution.
@@ III.1 Noise level without pump
Figure 2: a) Fluctuation of the electron diffraction pattern of Graphite sample in the absence of a pump as a function of time for an exposure time of one second, with a flux of $3.6\cdot 10^{6}$ electrons/sec. b) Spectral amplitude of the electron fluctuation as a function of the frequency, which corresponds to the Fourier transform of Fig 2a. c) Acquired count without electron beam as a function of the integration time due to cosmic ray having a flux of 3.1 events/sec from a linear fit. d) Variation of the noise level as a function of the cumulative signal $N$ for different configurations. The experimental curve is fitted with equation 2.
We acquired a series of diffraction patterns on a 50 nm thick Graphite single crystal. Data were recorded as a function of pump-probe delay time with the shot-to-shot acquisition system. For each acquisition window $(i)$, two diffraction patterns corresponding to $I_{\text{on}}$ and $I_{\text{off}}$ are collected while blocking the photoexcitation pulse before the sample. We sum the electron count over all the pixels to compute the total electron flux. Fig. 2a) shows the fluctuation of the normalized scattered electron flux for a single image, $\sum_{\text{px}}I_{\text{on}}$, and for the shot-to-shot acquisition, $\sum_{\text{px}}I_{\text{on}}/\sum_{\text{px}}I_{\text{off}}$. We see that the fluctuation amplitude decreased by more than one order of magnitude with the shot-to-shot acquisition than the conventional method.
Fig. 2b) shows the frequency spectrum of the normalized signal. The relatively flat spectrum suggests that the noise is essentially composed of white noise in both acquisition modes. The shot-to-shot acquisition reduces the noise level at every time regime by a factor of 40 in low frequency and 30 in high frequency.
@@
The source noise originates from fluctuations in the electron generation and the laser source which scales as $\sigma_{\text{source}}=\alpha_{s}\sqrt{f}\sqrt{N}$, with the electron flux $f$ and the RMS magnitude $\alpha_{s}$.
The white noise corresponds to the jitter and thermal effect of the optical and electron pulse generation scaling as $\sigma_{\text{white}}=\alpha_{w}N$.
The detector noise level defined as $\sigma_{\text{readout}}$ and $\sigma_{\text{int}}$ depends on the number of counts acquired at instantaneous and different integration times respectively when the electron beam is absent. The single electron detector doesn’t have thermal noise like the CCD detector, then $\sigma_{\text{readout}}=0$. This feature is crucial for utilizing the shot-to-shot acquisition method, as it prevents noise accumulation in each acquisition window. The main source of the integration noise $\sigma_{\text{int}}$ is the cosmic rays detected by the Quadro. However, as shown in Fig. 2c), the flux of cosmic rays of 3.1 events/sec is insignificant compared to the electron flux usually set to $\sim 10^{6}$ electrons/sec. As described in Sec. II.3, in the shot-to-shot mode, the exposure window is 25 $\mu s$ so over one second the effective exposure time of the detector is 0.5 sec, reducing by a factor of two the integration noise level. The gain noise, $\sigma_{\text{gain}}$, originates from converting a collision event in the silicon detector to a written bit. As explained in Appendix A, this noise is negligible as long as we don’t reach a saturation point of 1 electron/pixel/pulse.
@@
with $\alpha=(1+\alpha_{s}\sqrt{f})$. Fig. 2d) shows the noise level as a function of acquired electrons for single images and shot-to-shot mode for a flux of $f=3.6\cdot 10^{6}$ electrons/sec. It shows a source noise of $\alpha_{s}=10^{-2}$ Hz-1/2 and a white noise of $\alpha_{w}=1.7\cdot 10^{-2}$ for the single image whereas the shot-to-shot acquisition mode noise level reaches the shot noise with a source noise of $\alpha_{s}=1.6\cdot 10^{-4}$ Hz-1/2 and a white noise, $\alpha_{w}=1.5\cdot 10^{-16}$, considered negligible.

### v2

@@ Abstract
We demonstrate a novel shot-to-shot acquisition method for optical pump - keV electron energy probe in ultrafast scattering experiments. We integrate a phase-locked acquisition scheme at a repetition rate of 20kHz in a conventional ultrafast electron diffraction (UED) setup. We proceed to a full characterization of the noise level in different configurations and for realistic scenarios. The shot-to-shot acquisition improves the signal-to-noise ratio (SNR) by one order of magnitude and can be readily implemented in other high-repetition-rate electron diffraction and spectroscopy setups.
@@ I Introduction
Ultrafast electron diffraction, microscopy, and energy loss spectroscopy are widely used to investigate the properties of materials out-of-equilibrium using pump-probe methods. Transient signals recorded by these techniques correspond to a variation in the number of scattered electrons at specific positions in the momentum, real, or energy space, which enables studies in the temporal domain of lattice dynamics Carbone et al. (2008); Chatelain et al. (2014), magnetic ordering Rubiano da Silva et al. (2018a); Berruto et al. (2018); Rubiano da Silva et al. (2018b) and demagnetization Tauchert et al. (2022), plasmon dynamics Piazza et al. (2015), charge density waves coherent response and melting Cheng et al. (2024), and to resolve changes in the phonon population with momentum resolutionRené De Cotret et al. (2019); Stern et al. (2018); Waldecker et al. (2017). The amplitude of transient signals increases with the photoexcitation energy density and is typically on the order of a few percent for light excitation in the range of mJ/cm2. Advanced materials science focuses on broken-symmetry ground states, where a delicate microscopic balance of electron, spin, lattice, and orbital degrees of freedom is achieved. In electron diffraction, elastically scattered electrons provide measurements of lattice order, while inelastically scattered electrons contain crucial information on the momentum-dependent interaction between collective modes van der Veen, Penfold, and Zewail ; Stern et al. (2018). For instance, phonon populations alter thermal diffuse scattering, with a typical total cross-section of $10^{-4}$. Other quasiparticles, such as magnons, are also present in the diffuse inelastic scattering but represent only $10^{-6}$ of the total electron flux Lyon et al. (2021); Kepaptsoglou et al. (2025). Time-resolved measurements of such inelastic signals present a significant challenge for current experimental setups, which must improve detection sensitivity to capture small photo-induced variations, particularly in weak perturbation regimes that maintain the integrity of the ground state.
@@
To improve the signal-to-noise ratio (SNR), a first possibility is to develop a brighter electron source ($>10^{5}$ electrons/pulse). However, this approach introduces space-charge-induced temporal broadening of the electron pulse, which significantly degrades the time resolution Siwick et al. (2002); Wang et al. (2009). Temporal compression of the electron bunch can be achieved using a synchronized radio frequency cavity Mancini et al. (2012); Chatelain et al. (2012), or a terahertz optical pulse Ehberger et al. (2019). However, special care must be taken to avoid jitter and long-term drifts Otto et al. (2017). A second approach is to minimize the noise of the electron source while working at low electron flux ($<10^{4}$ electrons/pulse). To mitigate space charge broadening, the electron source is placed as close as possible to the sample Waldecker, Bertoni, and Ernstorfer (2015). Combined with a high repetition rate laser source and fast acquisition systemFreelon et al. (2023); Aidelsburger et al. (2010), the SNR is increased while maintaining good beam coherence. A possible limitation of this method is the sample re-equilibration time, which can become greater than the inter-pulse separation ($<1\mu$s) Vidas et al. (2020).
In this work, we present an ultrafast electron scattering apparatus based on a 20 kHz repetition rate source operating at low flux, with fewer than one thousand electrons per pulse. In this limit, the quality of the acquisition system is essential to improve the SNR. We employ a direct electron detector [^1] Quadro, Dectris, which records high SNR electron diffraction patterns McMullan, Faruqi, and Henderson (2016) thanks to a combination of high detective quantum efficiency (DQE) (0.9 for direct detection, as compared to 0.2 for a typical charge-coupled device (CCD)) Levin (2021) and low acquisition noise. Another advantage of such a direct detection system is its capability to acquire data at a high rate, synchronized with the laser repetition rate Duncan et al. (2023). This enables the collection of pump-probe data by chopping the pump pulse train at 10 kHz and normalizing the data using unpumped acquisition frames. This method, well-established in ultrafast optical spectroscopy due to the availability of fast light detectors, is widely used to achieve shot-noise-limited data Lang (2018). In electron scattering experiments, where the number of events per pixel and per pulse is low, the method remains viable due to the absence of the detector’s dark count contribution. We demonstrate pump-probe data with a noise level reaching $10^{-4}$ in a few hours of acquisition time, with sub-picosecond time resolution.
@@
The paper is organized as follows. In section II, we describe the setup, the acquisition system, and the working principles of the synchronization to achieve shot-to-shot acquisition. Section III presents the characterization of the noise level. In section IV, we will present UED data acquired in a realistic scenario by photo-exciting graphite samples with an 800 nm pump.
@@ II Design of the shot-to-shot UED
Figure 1: Scheme of the UED setup in EPFL composed of the optical table, the rack, and the computer. The optical table supports the pulsed laser, the optical beam path divided into the pump and the probe arms, the vacuum sample chamber, the photodiode, and the electron detectors. The rack accommodates the synchronization system, which is phase-locked to the laser source at 20 kHz. The data gathered by the electron and photodiode detectors are sent to the computer to be sorted between pumped and unpumped diffraction patterns.
@@ II.1 Optical setup and electron diffraction beamline
The ultrafast electron diffraction (UED) setup and the data acquisition scheme are schematized in Fig. 1. The current setup is a modification of the setup detailed in Ref. Mancini et al. (2012). A cryogenically cooled Ti:Sapphire laser [^2] Wyvern, KM Labs provides 60 fs pulses at a central wavelength of 780 nm at a repetition rate of 20 kHz with an output power higher than 10 W (pulse energy > 0.5 mJ). To avoid long-term drifts, the laser beam pointing is stabilized by a home-built stabilization system based on a pair of piezoelectric-actuated mirrors and CMOS cameras. The beam is divided into two arms with a beam splitter (BS1). The pump beam goes through a mechanical delay line to vary the delay between the probe pulse and the pump pulse. It then passed through a mechanical chopper, synchronized with the laser source, in order to halve the pulse repetition rate to 10 kHz. The last mirror (BS2) sends the pump into the chamber and is used to find spatial overlap between the pump beam and the electron probe. BS2 also acts as a beam splitter, transmitting a small fraction of the pump towards a photodiode.
@@
In the probe arm, third harmonic generation (THG) is used to generate a UV light beam (260 nm) which back-illuminates a silver-coated sapphire photo-cathode. The UV probe pulse energy exceeds the work function of the silver ($P_{\text{UV}}=4.76\text{ eV}>W_{\text{Ag}}=4.26$ eV), thus generating a pulse of electrons by single-photon photoemission. In Appendix A, we find the optimal condition for the photoemission when the measured quantum efficiency is maximum at $10^{-8}$ electrons/photon. It corresponds to an acquired flux between 1 and 1000 electrons/pulse photoemitted at UV powers from 1 $\mu$W and 1 mW, respectively. These pulses are then accelerated to 40 keV over a distance of 3 mm, corresponding to an accelerating electric field of 13.3 MV/m. The off-axis photoemitted electrons are filtered out using a 100 $\mu$m pinhole after the anode, providing the low beam emittance needed for coherent electron scattering with high reciprocal resolution. A solenoidal magnetic lens with approximately 3000 windings, each carrying a current of 0.62 A, focuses the electron beam onto the detector. Operating at 17.3 W, the lens achieves a focal length of 58 mm. At the sample holder position, 15 cm away from the electron source, the electron beam waist has a diameter of approximately 200 $\mu$m. For an electron flux of nearly 1000 electrons/pulse on the detector, we obtain a time resolution of around 1 ps.
@@
In transmission geometry, the sample holder hosts five standard 3 mm TEM grids. The holder is mounted on a 4-axis open-cycle cryo-manipulator that can achieve a temperature of 4.2K with liquid helium cooling. The UED setup can also work in reflection geometry to probe the surface lattice dynamics with five axes of freedom using another sample holder Pennacchio et al. (2017).
@@
The vacuum chamber has a base pressure of $10^{-8}$ mbar and accommodates a mirror that reflects the pump beams onto the sample at an angle of a few degrees off from the electron beam.
@@
The electron beam travels a distance of 57 cm after the sample position. At this distance and with a chip detection size of $38.4\times 38.4$ mm2, a reciprocal space area of $16\times 16$-2 is covered at 40 keV.
@@ II.2 High speed lock-in direct electron detection
We use this detector feature to acquire ultrafast electron diffraction by chopping the pump pulses at half the laser repetition rate (10 kHz). The two output images correspond to the excited, $I_{\text{on}}$, and non-excited, $I_{\text{off}}$, diffraction patterns of the studied sample. Due to the high shot-to-shot correlation of the laser source, this considerably reduces the noise in the pump-probe signal, as will be discussed in Sec. III. We now describe the synchronization system in more detail.
@@ II.3 Synchronization system
As shown in the left panel of Fig. 1, a homemade delay generator constantly receives the 20 kHz laser signal from the laser source. This trigger is relayed into four independently delayed channels (A-D). Channel A is frequency-divided to 10 kHz and transmitted to the chopper controller. Channel B triggers a fast photodiode digitizer  [^3] GlazPD, Synertronic design. This signal in this channel is gated by the main detector (EnOUT), which is a replica of the one received (EnIN). In this way, the Quadro detector operates as the master and the photodiode as the slave, ensuring that the same pulses originating from the laser are detected. The electron detector acquisition window is controlled by channel D, with a 20 kHz repetition rate of a 25 $\mu$s wide square signal. The acquisition window from channel D is inhibited by channel C, acting like a switch controlled by the main computer to start the acquisition. This synchronization system enables robust acquisition of every electron bunch at 20 kHz and can be readily adapted to other repetition rates. However, operation at higher frequencies would require a different chopping mechanism. From the main computer, the number of exposition windows as well as other detector settings can be monitored.
The main computer accesses the two 2D datasets acquired by the electron detector, $I_{\text{A}}$ and $I_{\text{B}}$, which can be sorted into $I_{\text{on}}$ and $I_{\text{off}}$, thanks to the photodiode readout. A Python-based user interface Usai and Claude (2024) displays this data and the difference between $I_{\text{on}}-I_{\text{off}}$, allowing for a fast and accurate time and spatial overlap.
@@ III Characterization
Given the low shot-to-shot fluctuation of the laser source, this acquisition method greatly improves the SNR of ultrafast experiments. We characterize this enhancement by disentangling the noise sources for conventional and shot-to-shot acquisition techniques with and without the photoexcitation pulse.
@@ III.1 Noise level without pump
Figure 2: a) Fluctuation of the electron diffraction pattern of graphite sample in the absence of a pump as a function of time for an exposure time of one second, with a flux of $3.6\cdot 10^{6}$ electrons/sec. b) Spectral amplitude of the electron fluctuation as a function of the frequency, which corresponds to the Fourier transform of Fig. 2a. c) Acquired count without electron beam as a function of the integration time due to cosmic ray having a flux of 3.1 events/sec from a linear fit. d) Variation of the noise level as a function of the cumulative signal $N$ for different configurations. The experimental curve is fitted with equation 2.
We acquired a series of diffraction patterns on a 50 nm thick graphite single crystal. Data were recorded as a function of pump-probe delay time with the shot-to-shot acquisition system. For each acquisition window $(i)$, two diffraction patterns corresponding to $I_{\text{on}}$ and $I_{\text{off}}$ are collected while blocking the photoexcitation pulse before the sample. We sum the electron count over all the pixels to compute the total electron flux. Fig. 2a) shows the fluctuation of the normalized scattered electron flux for a single image, $\sum_{\text{px}}I_{\text{on}}$, and for the shot-to-shot acquisition, $\sum_{\text{px}}I_{\text{on}}/\sum_{\text{px}}I_{\text{off}}$. We observe that the fluctuation amplitude is lower by more than one order of magnitude with the shot-to-shot acquisition than with the conventional method.
Fig. 2b) shows the frequency spectrum of the normalized signal. The shot-to-shot acquisition reduces the noise level at every time regime by a factor of 35 in low frequency and 20 in high frequency.
@@
The source noise originates from fluctuations in the electron generation and the laser source, which scales as $\sigma_{\text{source}}=\alpha_{s}\sqrt{f}\sqrt{N}$, with the electron flux $f$ and the RMS magnitude $\alpha_{s}$.
The white noise corresponds to the jitter and thermal effect of the optical and electron pulse generation, scaling as $\sigma_{\text{white}}=\alpha_{w}N$.
The detector noise level, defined as $\sigma_{\text{readout}}$ and $\sigma_{\text{int}}$, depends on the number of counts acquired at instantaneous and different integration times, respectively, when the electron beam is absent. As the single electron detector doesn’t have thermal noise like a CCD detector, $\sigma_{\text{readout}}=0$. This feature is crucial for utilizing the shot-to-shot acquisition method, as it prevents noise accumulation in each acquisition window. The main source of the integration noise $\sigma_{\text{int}}$ is the cosmic rays detected by the single electron detector. However, as shown in Fig. 2c), the flux of cosmic rays of 3.1 events/sec is insignificant compared to the electron flux usually set to $\sim 10^{6}$ electrons/sec. As described in Sec. II.3, in the shot-to-shot mode, the exposure window is 25 $\mu s$ per pulse, so over one second, the effective exposure time of the detector is 0.5 s, reducing the integration noise level by a factor of two. The gain noise, $\sigma_{\text{gain}}$, originates from converting a collision event in the silicon detector to a written bit. As explained in Appendix A, this noise is negligible as long as we don’t reach a saturation point of 1 electron/pixel/pulse.
@@
with $\alpha=(1+\alpha_{s}\sqrt{f})$. Fig. 2d) shows the noise level as a function of acquired electrons for single images and shot-to-shot mode for a flux of $f=3.6\cdot 10^{6}$ electrons/sec. It shows a source noise of $\alpha_{s}=10^{-2}$ Hz-1/2 and a white noise of $\alpha_{w}=1.7\cdot 10^{-2}$ for the single image, whereas the noise level of the shot-to-shot acquisition mode reaches the shot noise limit with a source noise of $\alpha_{s}=1.6\cdot 10^{-4}$ Hz-1/2 and a white noise, $\alpha_{w}=1.5\cdot 10^{-16}$, considered negligible.